Oxford Materials IT Committee Meeting

MICHAELMAS TERM 2007
A meeting of the Department of Materials IT Committee was held on Tuesday 30th October 2007.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters arising from the Minutes of the Meeting of the last meeting
  • 3 IT/AV – Matters for Discussion (attached)
  • 4 IT/AV – Head of IT Report (attached)
  • 5 IT/AV expenditure - matters for report (PJW)
  • 6 AOB
  • 7 Date and time of next meeting: Tuesday 29th January 2008 at 2pm in HR Meeting Room

ITEM 2. MATTERS ARISING FROM THE MEETING

See minutes below.

ITEM 3: IT/AV MATTERS FOR DISCUSSION

No additional items tabled for discussion.

ITEM 4: IT/AV HEAD OF IT REPORT

Admin Network
There was a catastrophic hardware failure on Materials2 in July followed by further instability under high loads. At the end of August it was decided to rush forward existing plans for migrating all users from Materials2 to Materials0. Some migrations have taken place already and further migration of users is on-going.

Teaching Laboratory
All the computers in the computer room have been migrated from Materials2 to Materials0 and all undergraduate accounts have been migrated. The laptops and microscope pc’s remain to be migrated. One pc had it’s case replaced due to a broken power switch.

Lecture theatres / meeting rooms
Over the summer PJW and AOT produced a plan to upgrade the lecture theatres and meeting rooms using Roberts money. A questionnaire was circulated to seniorstaff@materials and the feedback compiled. From the feedback it was determined that there was interest in most of the proposed upgrades, therefore after minor revisions three companies were approached to obtain quotes for the work in each of the four rooms. The upgrade of HR lecture theatre and HR meeting room took place just before the start of term however problems encountered meant that the equipment was not fully operational until mid term and instruction sheets and staff training still remain to be done. The BR and ETB committee rooms will be upgraded in November/December and the staff training organised over the Christmas vacation ready for next term.

Digital Darkroom
The desktop pc’s are about to be upgraded (with slightly newer trickle-down pc’s) when they are moved from Materials2 to Materials0.

Library
The desktop pc’s are about to be upgraded (with slightly newer trickle-down pc’s) when they are moved from Materials2 to Materials0.

MML
The new quadcore cluster was expanded from 9nodes to 17 nodes. The plantroom refurbishment project was successfully completed integrating with the full power outage of the building associated with re-routing the power for 12/13PR. The department now has a server-room suitable for housing the MML clusters, the department servers, the electron microscopy servers and workstations and miscellaneous other equipment. See attached diagram and photos.

EM
Chao Lin was interviewed and appointed by AIK (with PJW). He has taken over IT support of for AIK group activities. So far there has been little integration with department IT support, however it is still hoped to train Chao to perform so core IT tasks over and above his group activities. Plans for refurbishment of the EMIT lab are close to being finalized. The planned upgrade of the HR<->HB link to gigabit encountered problems and needs to be readdressed. RCD has requested re-networking of the old STEM room which is being outsourced to Monards.

Website
During Trinity term PJW completed an initial review of the possible content management systems (CMS), then selected several contenders (drupal,joomla,plone,cmsmadesimple) and created draft installations and carried out usability/functionality tests. Plone was ruled out as being python rather than php. Drupal and Joomla although initially interesting proved less user friendly than CMS Made Simple.
In July PJW ported the QIPIRC website into CMS Made Simple as a self-training experience (2days). [ Note that the QIPIRC are still updating the content of their website so the new site has not actually gone live yet.]
In August PJW created an initial functional template similar to the existing department website then in September ported the whole of the www.materials.ox.ac.uk website (~300pages) into the new CMS which then went live on 5th October right at the busiest time of year!
During Michelmas Term it is intended to arrange consultation meetings about the design and structure of website leading to significant further redevelopment of the look of the website. This should also review the content delivery to target audiences (surfers, prospective students, collaborators, existing students and staff, industry etc.) In addition further technological development database information are intended for presenting research projects, new projects available, publications, research highlights, department news, a department diary, private content for internal view only etc.
There are also on-going activities in relation to several other websites such as www-solar.materials.ox.ac.uk, outreach.materials.ox.ac.uk, matsoc.materials.ox.ac.uk, etc.

Publications
The new handbooks were collated and printed ready for the new academic year. The Research in Progress handbook has been postponed due to the changes in the website, but it is intended to complete RIP2007 before Christmas.

Network
New Edgeiron switches were installed in PR, BR, BB and HB, and the links to ETB, BR, BB were successfully upgraded to gigabit. HB encountered problems but will be upgraded to gigabit soon.

Servers
The new server room was commissioned in July and most servers have successfully migrated into the new room. The Materials2 fileserver failed and motherboard and case/psu were replaced but there have been ongoing problems of stability under high load so the planned migration to Materials0 has been rushed forward and AMC has been working hard to develop the new system and migrate users.
The management server oums-nomad had a memory upgrade and had both mirrored disks replaced due to hardware failures. The windows update server oums-wsus had a memory upgrade.
The old windows update server oums-sus was decommissioned and the network monitoring utility Transcend was migrated to oums-alex which backs up the Secure volume locally. The MML raid array was replaced and the old raid array is ready for decommissioning. The Emgroup1 raid array only has 10 users which can be migrated onto the MML raid array later this term prior to decommissioning. A new server was commissioned (painfully) for OMCS to use ACT!2007.

Phones
Six new VOIP phones were purchased and installed at Begbroke in the space vacated by OGT. Numerous problems with the VOIP phones occurred over the summer – hardware was replaced at the exchange, configuration changes were made at the exchange. Some problems were tracked to a local network mismatch. There are still intermittent reports of problems and several outstanding hardware repair requests.

Security
Several instances of possible p2p activity were investigated but no action taken formally. In August there was a spate of viruses in the PSG group and the EM group. These were quite nasty viruses which seemed to propagate/reinfect via memory sticks. Infected machines were difficult to analyse due to Sophos only detecting secondary evidence of infection. Eventually around 10 machines were reinstalled.

Disposal
The new WEEE rules came into force in July. The email request to dispose of equipment before the deadline did not achieve the desired effect, but now there is a backlog of equipment needing disposal urgently (before the next safety inspection!).

Software
Since the Origin became available under a departmental-site-license we have installed Origin on 56 desktops and 30 laptops.
Over the summer, Mike Giles at OSC has been gathering support across the MPLS Division to add another licensing method for Matlab such that we will be able to roll-out basic Matlab installations onto all undergraduate and postrgraduate student personally owned computers (which will work off the network). The department has committed to an initial contribution of £600 this year in support of this and the license should be available by Christmas.


MINUTES

A meeting of the Department of Materials IT Committee was held on Tuesday 30th October 2007.
The following people attended: APD, PJW, KS, AHM, JAMS and CJS.

1. Apologies for Absence and Membership of the Committee

Apologies were received from Ralf Drautz due to examiners meeting and Alex Cock due to door-entry system engineers being on-site.
CJS noted that his membership of the committee expired during the summer therefore he duly resigned. The committee thanked him for his input and for representing users at Begbroke.
The committee discussed several possible new members of the committee in terms of their potential for representing the different groups of users around the department (e.g. Begbroke, EM, QIP, MML etc). PJW will take our recommendations to the Head of Department ready for next term.

2. Matters arising from the minutes

a) IT disaster recovery plans (Minute 2a)
The department now has a fit-for-purpose server room. The core network devices and the servers within that server room are connected to a 5kVA UPS (currently running at 9A load). This UPS will provide some (<2hr) cover in the event of disruption to power, however it is not intended to be able to cover an overnight or weekend outage.
Since many departmental servers are now co-located, it is important to ensure that data is backed-up to a different location in the event of flood/fire in the server room.
An LTO4 tape drive (800Gb native / 1600Gb compressed) with a dedicated server has been purchased and is being commissioned by KS. Once the tape drive is successfully operational as the disaster recovery backup of the MML and departmental fileservers, then the system will be assessed for it’s ability to provide additional backup or even archiving of data at a department level (alternative to the university HFS Archive system).
Catastrophic hardware failure of Materials2 occurred in July. Replacement hardware was sourced purchased and installed and the system was operational again in less than 48 hrs. This is a fair example of a real disaster recovery scenario. No data was lost. More significant has been the on-going instability of the server after the hardware replacement.
There was considerable discussion of the impact of the instability of Materials2 upon the core function of the department. PJW updated the committee about the current plans for migrating users.
CJS noted that the total failure of the Q drive (both disks) in mid-August left OMCS unable to access their data for 4 days, which was extremely inconvenient. CJS requested that procedures be changed to improve the response time for such events.
There is still a need for disaster recovery documentation to ensure that there is adequate provision and there is a clear structure for achieving efficient disaster recovery.

Action: PJW

b) Website (Minute 2b)
The outstanding proposed review of research group websites (since summer 2006) still has not happened however the developments in the department website now enable this to move forward during the next academic year.

Action: PJW

c) Copyright (Minute 2c)
APD/PJW chased staff for returns, to request copyright signs be placed beside all scanners and to remind staff about emailing URL's rather than actual copyrighted PDF's.

d ) Website (Minute 2d)
The outstanding request that the graduate admissions section of the website be updated has still not progressed since being discussed by PJW and JAMS however the developments in the department website now enable this to move forward during the next academic year.

Action: PJW/JAMS

e) Server provision and proposal for server room (Minute 2e)
The new server room was completed on schedule in July and became fully operational. (Further details in Head of IT report below)

f) Email repository (Minute 2f)
The proposed departmental email repository has been installed and is being used (currently only 8 users). The same server has also taken over running the public FTP service and the previous temporary server has been decommissioned. Both services have been advertised to users by email and via the website.

g) Usage of spare CPU capacity for Modelling (Minute 2g)
RD on behalf of the MML proposed that technical solutions be investigated for utilising spare CPU capacity around the department for running Modelling jobs. An initial aim could be to run jobs on the Teaching Laboratory computers at Weekends and if successful there may be scope for extending this to other departmental computers and other time frames. RD, KS and PJW to investigated using VMplayer to link into the MML queuing system but did not actually achieve a working system. OeRC are also proactively developing solutions in this area. Item carried forward.

Action: RD/KS/PJW

h) Video-conferencing (Minute 2h)
At the meeting of senior staff in July it was decided to discontinue the webcasting project as being too time consuming and too expensive in hardware costs for the benefit obtained. Over the summer Physics continued to investigate replacing the webcasting equipment with a more user-friendly system.

i) Department Website Re-design. (Minute 2i)
The first phase of redevelopment of the department website went live on 5th October. (see more details in the Head of IT report below).
There was considerable discussion about the recent upgrade to the department website. The slow speed of the website was noted, and PJW confirmed that we urgently need to move the site to a newer faster webserver.
The planned consultation meetings about the design and structure of website leading to significant further redevelopment of the look of the website need to be arranged so that the project moves forward more quickly.

Action: PJW/Kay Davies

j) Archive procedures. (Minute 3a)
Currently IT staff have several different archive issues under consideration: 1) archiving data off servers when people have left 2) implementing an archive procedure for OMCS that satisfies ISO9001. 3) Archiving FIM group data from the new atom probes 4) archiving procedure for collaborative storage areas.

Archive to HFS seems to be a good service but requests data “is of value to the University”.

Archvie procedure and service should vary according to disk space requirements.

Archive service should be chargeable to discourage “just in case” excessive archiving.

Group leaders are presumed to be responsible for archiving within their groups.

Questionnaire for group leaders would help clarify the current status and promote best practice.
No action over the summer. Item carried to next meeting.
KS offered to look into the specific requirements of OMCS for ISO9001.

Action: AMC/KS/PJW

k) Lecture theatre and meeting room equipment. (Minute 4a)
Over the summer PJW and AOT produced a plan to upgrade the lecture theatres and meeting rooms using Roberts money. A questionnaire was circulated to seniorstaff@materials and the feedback compiled. From the feedback it was determined that there was interest in most of the proposed upgrades, therefore after minor revisions three companies were approached to obtain quotes for the work in each of the four rooms. The upgrade of HR lecture theatre and HR meeting room took place just before the start of term, however problems encountered meant that the equipment was not fully operational until mid term and instruction sheets and staff training still remain to be done. The BR and ETB committee rooms will be upgraded in November/December and the staff training organised over the Christmas vacation ready for next term.

Action: PJW/AOT

3 Matters for Discussion

No additional items were tabled for discussion.

4. Head of IT Report

A written report was circulated with the agenda. Key points were highlighted.
JAMS noted that the Origin site-license seems to have been very useful.
KS have a brief verbal report on the recent changes to the network.

5. IT/AV Expenditure

No financial data was presented at the meeting.
Subsequent analysis of the 2006/7 end-of-year finances showed a total non-staff expenditure of around £54k however £13k has already been reimbursed from grants for equipment provided and a further £3k needed reclaiming so the actual expenditure was around £38k (and even that included the previous year’s server monitoring charge of £2.7k). This fell within the overall budget of £40k.
In the first three months of this financial year expenditure has been around 17.5-4.5=£13k.

6. AOB

a) Oxford Research Archive (ORA)
APD asked whether the members of the ITC had heard of the ORA. PJW commented that the ORA is a re-launch of a previously failed project to create an institutional archive. The main problem with the previous project had been lack of uptake primarily due to copyright restrictions and due to researchers needing to publish in high-impact journals in order to achieve maximum RAE funding. PJW noted that Prof. George Smith has been involved in this current ORA project and is probably better placed to comment on whether the project is likely to be of interest to members of department. PJW to approach GDWS.

Action: PJW

7. Date of Next Meeting

The next meeting will take place on Tuesday 29th January 2008 at 2pm in HR Meeting Room (room booked already)
The meeting closed at 3.30pm